# Covariate Adjusted Response Adaptive Design with Delayed Outcomes

### v1

@@ 2 Optimal delay-adjusted treatment allocation in CARA
Design objective 2. Optimal treatment allocation in design on failure reduction under a power constraint. This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger et al., 2001). In this context, a treatment is considered to fail when the potential outcome equals one. The optimal allocation thus seeks to assign treatments based on the solution to the following optimization problem:
$\min_{e_{1},\ldots,e_{T}}$ $\ \sum_{t=1}^{T}r_{t}\sum_{x\in\mathcal{X}}p(x)\Big{[}e_{t}(1|x)\left(1-\mu(x,1)\right)+\left(1-e_{t}(0|x)\right)\left(1-\mu(x,0)\right)\Big{]}$ (3)
s.t. $\ \delta\leq e_{t}(\cdot)\leq 1-\delta,\quad\text{for }t=1,\ldots,T$
$\ \mathsf{V}\left(e_{1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}p(x)\big{(}\tau(x,1,0)-\tau(1,0)\big{)}^{2}\leq C$
In this formulation, the objective function minimizes the total expected number of failures during the entire experiment. The term inside the summation accounts for the expected failures when assigning treatment 1 with probability $e_{t}(\cdot)$ and treatment 00 with probability $1-e_{t}(\cdot)$. Here, $\mu(x,a)$ represent the expected success probabilities under treatments $a$, given covariates $x$. The first feasibility constraints again ensure that the treatment assignment probabilities remain within the interval $[\delta,1-\delta]$. This design balances the ethical imperative to reduce adverse outcomes with statistical consideration to maintain sufficient power for detecting treatment effects.
Due to the delay mechanism’s influence on the variance lower bound, the optimal treatment allocation rules differ substantially from those derived in the existing literature. We shall show that the solutions to the optimization problems in Eq (2) and Eq (3), denoted as $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$, from a sequence of probabilities that vary across different stages. In comparison, the classical Neyman allocation rule for power maximization:
${e}_{t}^{\mathtt{Neyman}}(a|x)=\frac{\sigma(x,a)}{\sigma(x,1)+\sigma(x,0)},\quad a\in\{0,1\},\quad t=1,\ldots,T,$ (4)
and the optimal failure reduction rule derived in Rosenberger et al. (2001):
${e}_{t}^{\mathtt{FR}}(a|x)=\frac{\sqrt{\mu(x,a)}}{\sqrt{\mu(x,1)}+\sqrt{\mu(x,0)}},\quad a\in\{0,1\},\quad t=1,\ldots,T,$ (5)
@@
Figure 2: Comparison between classical optimal rule and delay-adjusted optimal rule for two design objectives in CARA: (A)-(B) power maximization and (C)-(D) reducing the total number of failures.
We illustrate this point more concretely in a simulation study, where model parameters are calibrated from our synthetic case study and are applied to the oracle optimization problems in Eq (2) and (3). We refer to the detailed data-generating process for this simulation design in Section 5. First, we solve for the oracle treatment allocation with the design objective of maximizing power, as defined in Eq (2). The resulting oracle treatment allocations for each Stage $t$ and the Neyman allocation are presented in Panels (A) and (B) for covariate strata $X=0$ and $X=1$, respectively. The corresponding values of the objective function in Eq (2) under the two different allocations are labeled as $\mathsf{V}_{\text{Delay}}$ and $\mathsf{V}_{\text{Neyman}}$. Second, we solve for the oracle treatment allocation with the design objective of minimizing the total expected number of failures, as in Eq (3). The oracle treatment allocation and the ethical allocation across Stage $t$ are shown in Panels (C) and (D) for the covariate strata $X=0$ and $X=1$, respectively. The percentage of failure under the two allocation strategies is labeled in the figures.
As illustrated in Figure 2(A), the optimal treatment allocation accounting for the response delay (i.e., the solution to Eq (2)) varies across stages and deviates significantly from Neyman allocation. Furthermore, the optimal treatment allocations result in an ATE estimator with reduced variance (e.g., $\mathsf{V}_{\text{Delay}}=1.62$ and $\mathsf{V}_{\text{Neyman}}=2.95$ for the strata with $X=0$). As illustrated in Figure 2(B), a similar observation applies to ethical designs, where the optimal treatment allocation tailored to the ethical design objective (i.e., the solution to Eq (3)) differs from the ethical design rule in Eq (5), and the delay-adjusted optimal allocation achieves a lower failure rate.
Before concluding this section, we note that under appropriate conditions, we will show that the optimal treatment allocation rules $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$ are unique (Theorem 4). This uniqueness lays the foundation for establishing the statistical guarantees of our proposed design. Moreover, it demonstrates the necessity of accounting for delay mechanisms to achieve the desired experimental design goals.
@@ 3 Fully forward-looking CARA with delayed outcomes
The results provided in the previous section assume prior knowledge of the delay mechanism, and therefore, the discussed optimal treatment allocation corresponds to an infeasible oracle setting. In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective. In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev et al., 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger et al., 2001). We also provide tools for valid statistical inference after the experiment is finished.
@@
As with any CARA designs, at the first stage of the experiment, we lack information about the participants, so treatments are assigned randomly with a pre-fixed probability. In subsequent stages, our proposed fully forward-looking CARA design incorporates two key components: (1) estimating past delay mechanisms from collected data while anticipating future delays, and (2) using these estimates to formulate a fully forward-looking optimization problem that determines optimal treatment assignments for the next stage. An overview of our proposed fully forward-looking CARA can be found in Figure 3. We introduce these two components in the following paragraphs.
To save notation, we consider a slightly simplified setting in this section that there are only two arms, that is, $a\in\mathcal{A}=\{0,1\}$. We remark that our design strategies can be easily extended to a more general multi-valued treatment setting, with the cost of additional notation. In Section 4, we will establish properties of our proposed design methods and the treatment effect estimators under the general setting, accommodating multiple treatment arms.
Figure 3: Forward looking CARA design illustration.
At the end of each stage $t\geq 2$, we estimate (and update) the observed delay probabilities as
$\hat{\rho}_{t}(d|x,a)=\sum_{\ell=0}^{d}\hat{\mathbb{P}}[D_{i}=\ell|X_{i}=x,A_{i}=a]=\sum_{\ell=0}^{d}\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}=\ell,S_{i}\leq t-\ell)}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,S_{i}\leq t-\ell)}},\quad d\leq t-1,$
@@
for $a=0,1$. Notice that the above estimator is only defined for $d\leq t-1$. To gain some intuition, consider a concrete example for $t=3$. At the end of the third stage, data collected from participants enrolled in the first stage will be informative about the delay mechanism $\rho(0)$, $\rho(1)$, and $\rho(2)$. Among those whose outcome information is still missing at the end of stage 3, it is not possible to tell whether such information will become available until the end of stage 4. Following a similar reasoning, participants enrolled in the second stage can help estimate $\rho(0)$ and $\rho(1)$. The main message is that at any stage $t$, the delay mechanism is only estimable up to $d\leq t-1$, and the remaining “tail features” are not estimable until more data is collected. To anticipate these future delay distributions, we propose the following approaches that reflect the experimenter’s expectations about future delays. For a conservative experimenter, we consider a simple extrapolation approach:
$\hat{\rho}_{t}(d|x,a)=\hat{\rho}_{t}(t-1|x,a),\quad d\geq t.$
This extrapolation essentially assumes that all future delay probabilities will remain at the last estimated value. For an optimistic experimenter, we consider an extrapolation that postulates all missing outcomes will be available after one more experimental stage:
$\hat{\rho}_{t}(d|x,a)=1,\quad d\geq t.$
@@
Lastly, a neutral experimenter might gradually adjust the estimated delay probabilities from the last observed value to immediate response by interpolating between $\hat{\rho}_{t}(t-1|x,a)$ and 1 for $d=t,\dots,T-1$. These estimation strategies enable experimenters to incorporate future delay probabilities into our design, accommodating different perspectives on how delays may evolve. An illustration of different delay mechanism estimation strategies is plotted in Figure 4. Although it is generally impossible to rank the three extrapolation methods, we show through Monte Carlo experiments that they all deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches (such as complete randomization). To avoid unnecessary length discussions, we focus on the conservative approach in our theoretical investigation and show that the optimized treatment allocation is consistent (Theorem 4).
Figure 4: Comparison of different experimenter perspectives on delays and their corresponding estimation on the delay mechanism.
@@
Next, to optimize treatment allocation for stage $t+1$ while accounting for response delays, we introduce a fully forward-looking estimation of the variance term $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ at the current stage $t$. Our proposed estimate incorporates past allocations up to Stage $t$, future allocations from stages $t+1$ to $T$, and both past and future delay mechanisms. It is designed to anticipate the impact of future allocations and outcome delays on the variance of the final treatment effect estimate after the experiment is finished. Our estimate is of the following form:
@@
$\hat{\mathsf{V}}^{\text{forward}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
$\qquad\quad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
In this expression, $\hat{p}_{t}(x)$ represents the estimated proportion of participants with covariate $x$ up to stage $t$, calculated as
$\hat{p}_{t}(x)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,S_{i}\leq t)}}{N_{t}},$
@@
where recall that $n_{\ell}$ is the number of participants enrolled at stage $\ell$, $N_{t}$ is the total number of participants up to stage $t$, and $N=N_{T}$. The term $\hat{r}_{\ell}=n_{\ell}/N$ denotes the proportion of participants enrolled at stage $\ell$. The estimated conditional variance of the outcome $Y(a)$ given covariate $x$ and treatment $a$, up to stage $t$, is denoted by $\hat{\sigma}_{t}^{2}(x,a)$ and is calculated as
$\hat{\sigma}_{t}^{2}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}(Y_{i}-\hat{\mu}_{t}(x,a))^{2}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}},\quad\hat{\mu}_{t}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}Y_{i}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}}.$
While the expressions may seem complicated, they are simply the sample mean and sample variance of the outcome variable, restricted to the subsample for group $x$ receiving treatment $a$.
@@
$(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))=\operatorname*{argmin}_{e_{t+1},\ldots,e_{T}}\left\{\hat{\mathsf{V}}^{\text{forward}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)):\ \delta\leq e_{\ell}(\cdot)\leq 1-\delta,\ \ell=t+1,\ldots,T\right\}.$ (6)
We then set $\hat{e}^{*}_{t+1}(\cdot)=\tilde{e}_{t+1}(\cdot)$.
@@
$\hat{\mathsf{P}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)=$ $\sum_{\ell=1}^{t}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\left[\hat{e}^{*}_{\ell}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+\hat{e}^{*}_{\ell}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\right]$
$+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\left[e_{t}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+e_{t}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\right].$
@@
$\min_{e_{t+1},\ldots,e_{T}}$ $\ \hat{\mathsf{P}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)$ (7)
subject to $\ \delta\leq e_{\ell}(\cdot)\leq 1-\delta,\quad\text{for }\ell=t+1,\ldots,T$
$\ \hat{\mathsf{V}}^{\text{forward}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\big{(}\hat{\tau}_{t}(x)-\hat{\tau}_{t}\big{)}^{2}\leq C,$
where $\hat{\tau}_{t}(x)=\hat{\mu}_{t}(x,1)-\hat{\mu}_{t}(x,0)$ and $\hat{\tau}_{t}=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\hat{\tau}_{t}(x)$. We then again set $\hat{e}^{*}_{t+1}(\cdot)=\tilde{e}_{t+1}(\cdot)$.
To conclude our design strategy, as both optimization problems in Eqs  (6) and (7) are fully forward-looking, our proposed design strategies aim to proactively adjust the allocation strategies in future stages, ensuring that the data collected accounts for the future impacts due to outcome delays. This approach contrasts with classical CARA designs that optimize trial objectives solely using historical information. At the end of the last stage $T$, the average treatment effect estimator is constructed with
@@
$\hat{\mathsf{V}}_{T}$ $=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\Bigg{[}\frac{\hat{\sigma}_{T}^{2}(x,1)}{\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=1,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}}+\frac{\hat{\sigma}_{T}^{2}(x,0)}{\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=0,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}}+\big{(}\hat{\tau}_{T}(x)-\hat{\tau}\big{)}^{2}\Bigg{]}.$
In the following sections, we will provide theoretical guarantees and simulation results to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.

### v2

@@ 2 Optimal delay-adjusted treatment allocation in CARA
Design objective 2: Optimal treatment allocation for failure reduction subject to a power constraint. This design aims to minimize the expected number of failures while satisfying a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001). In this context, a treatment is considered to fail when the potential outcome equals one. The optimal allocation thus seeks to assign treatments based on the solution to the following optimization problem:
$\min_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\ \sum_{t=1}^{T}r_{t}\sum_{x\in\mathcal{X}}p(x)\Big{[}e_{t}(x)\left(1-\mu(x,1)\right)+\left(1-e_{t}(x)\right)\left(1-\mu(x,0)\right)\Big{]}\quad$ (4)
$\qquad\qquad\text{s.t.}\quad\ \delta\leq e_{t}(\cdot)\leq 1-\delta,\quad\ \mathsf{V}\left(e_{1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}p(x)\big{(}\tau(x)-\tau\big{)}^{2}\leq C.$
In this formulation, the objective function minimizes the total expected number of failures during the entire experiment. The term inside the summation accounts for the expected failures when assigning treatment 1 with probability $e_{t}(x)$ and treatment 0 with probability $1-e_{t}(x)$. Here, $\mu(x,a)$ represent the expected success probabilities under treatments $a$, given covariates $x$. The first feasibility constraints again ensure that the treatment assignment probabilities remain within the range $\delta,1-\delta]$. This design balances the ethical imperative to reduce adverse outcomes with statistical consideration to maintain sufficient power for detecting treatment effects.
Due to the delay mechanism’s influence on the variance lower bound, the optimal treatment allocation rules differ substantially from those derived in the existing literature. We shall show that the solutions to the optimization problems in (3) and (4), denoted as $e_{1}^{*}(\cdot),\ldots,e_{T}^{*}(\cdot)$, form a sequence of probabilities that vary across different stages. In comparison, the classical Neyman allocation rule for power maximization:
$e_{t}^{\mathtt{Neyman}}(a|x)=\frac{\sigma(x,a)}{\sigma(x,1)+\sigma(x,0)},\quad\text{for }t=1,\ldots,T,$ (5)
and the optimal failure reduction rule of Rosenberger, Stallard, Ivanova, Harper and Ricks (2001):
$e_{t}^{\mathtt{FR}}(a|x)=\frac{\sqrt{\mu(x,a)}}{\sqrt{\mu(x,1)}+\sqrt{\mu(x,0)}},\quad\text{for }t=1,\ldots,T,$ (6)
@@
Figure 2: Optimal oracle allocation (accounting for outcome delays) and the Neyman allocation for the power maximization objective.
We illustrate this point more concretely by computing the optimal oracle treatment allocation in the example introduced in Section 1, and compare it with classical treatment assignment methods such as the Neyman allocation. The covariate strata we consider are formed by two covariates, the biological sex and the WHO clinical stages of HIV. See Section 5 for additional details on the model parameters. In Figure 2, we present the oracle allocations for each stage and each covariate stratum corresponding to the power maximization objective defined in (3), as well as the Neyman allocation. Note that, since the classical Neyman allocation ignores the delayed outcome issue, it results in a constant treatment probability across the experimental stages, whereas the oracle allocation varies significantly across both stages and covariate strata. This highlights the importance of allowing the delay mechanism to be both arm- and covariate-specific.
We also report the objective function value for each design method in the figure. As can be seen, the efficiency gain in the oracle allocation can be substantial: more than 10% in the first covariate stratum and about 8% in the last case. Once again, this numerical exercise demonstrates the potential gains from leveraging the delay mechanism in a CARA framework.
@@ 3 Fully forward-looking CARA with delayed outcomes
The results provided in the previous section assume prior knowledge of the delay mechanism, and therefore, the discussed optimal treatment allocation corresponds to an infeasible oracle setting. In practice, the delay mechanism is rarely known and must be estimated as data are sequentially accumulated to optimize the experimental objective. In this section, we propose two delay-informed practical CARA designs tailored to two previously discussed experimental goals: improving statistical power (Tymofyeyev, Rosenberger and Hu, 2007) and minimizing the expected number of failures while meeting a minimum power requirement for binary responses (Rosenberger, Stallard, Ivanova, Harper and Ricks, 2001). We also provide tools for valid statistical inference after the experiment is finished.
@@
As with any CARA design, we use a pre-specified and fixed treatment assignment rule for the first stage due to the lack of a priori knowledge about the covariate and outcome distributions as well as the delay mechanism. In subsequent stages, our proposed fully forward-looking CARA design incorporates two key components: (i) estimating aspects of the delay mechanism from accrued data, while employing extrapolation to assess the tail behavior of the delay distribution; and (ii) using these estimates to formulate a feasible, fully forward-looking optimization problem that determines the optimal treatment assignments for the next stage.
We first discuss the estimation of the delay mechanism. Assuming stage $t$ of the experiment has concluded, we first update our delay distribution estimator via
$\hat{\rho}_{t}(d|x,a)=\sum_{\ell=0}^{d}\hat{\mathbb{P}}[D_{i}=\ell|X_{i}=x,A_{i}=a]=\sum_{\ell=0}^{d}\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}=\ell,S_{i}\leq t-\ell)}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,S_{i}\leq t-\ell)}},\quad d\leq t-1,$ (7)
@@
for $a=0,1$. Notice that the above estimator is only defined for $d\leq t-1$. To gain some intuition, consider a concrete example for $t=3$. At the end of the third stage, data collected from participants enrolled in the first stage will be informative about the delay mechanism $\rho(0|\cdot)$, $\rho(1|\cdot)$, and $\rho(2|\cdot)$. Among those whose outcome information is still missing at the end of stage 3, it is not possible to tell whether such information will become available until the end of stage 4. Following a similar reasoning, participants enrolled in the second stage can help estimate $\rho(0|\cdot)$ and $\rho(1|\cdot)$. The main message is that at any stage $t$, the delay mechanism is only estimable up to $d\leq t-1$, and the remaining “tail features” are not estimable until more data is collected.
To anticipate these future delay distributions, we propose the following approaches that reflect the experimenter’s expectations about future delays. For a conservative experimenter, we consider a simple extrapolation approach:
$\hat{\rho}_{t}(d|x,a)=\hat{\rho}_{t}(t-1|x,a),\quad d\geq t.$ (8)
This extrapolation essentially assumes that all future delay probabilities will remain at the last estimated value. As we discuss below in Section 4.3, the conservative extrapolation can be understood from a minimax perspective, as we will be sequentially optimizing the objective function under the worst-case scenario of the delay mechanism. For this reason, we adopt this approach in most of our numerical experiments and theoretical investigations.
@@
For an optimistic experimenter, we consider an extrapolation that postulates all missing outcomes will be available after one more experimental stage: $\hat{\rho}_{t}(d|x,a)=1,d\geq t$. Lastly, a neutral experimenter might gradually adjust the estimated delay probabilities from the last observed value to immediate response by interpolating between $\hat{\rho}_{t}(t-1|x,a)$ and $1$ for $d=t,\dots,T-1$. These strategies enable experimenters to incorporate future delay probabilities into our design, accommodating different perspectives on how delays may evolve. Although it is generally impossible to rank the three extrapolation methods, we show through Monte Carlo experiments that they all deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches (such as complete randomization).
@@
Next, to optimize treatment allocation for stage $t+1$ while accounting for response delays, we introduce a fully forward-looking estimation of the variance term $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ at the current stage $t$. Our proposed estimate incorporates past allocations up to Stage $t$, future allocations from stages $t+1$ to $T$, and the estimated delay mechanisms. Our method is designed to anticipate the impact of future allocations and outcome delays on the variance of the final treatment effect estimate after the experiment concludes:
@@
$\hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Bigg{[}\frac{\hat{\sigma}_{t}^{2}(x,1)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)\hat{e}^{*}_{\ell}(1|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,1)e_{\ell}(x)}$
$\qquad\qquad+\frac{\hat{\sigma}_{t}^{2}(x,0)}{\sum\limits_{\ell=1}^{t}\hat{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)\hat{e}^{*}_{\ell}(0|x)+\sum\limits_{\ell=t+1}^{T}{r}_{\ell}\hat{\rho}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
@@
Here, $\hat{p}_{t}(x)=N_{t}^{-1}\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,S_{i}\leq t)}$, which is the estimated proportion of participants with covariate $x$ up to stage $t$. The term $\hat{r}_{\ell}=n_{\ell}/N$ denotes the proportion of participants enrolled at stage $\ell$. The estimated conditional mean and variance of the outcome $Y_{i}(a)$ are
$\hat{\sigma}_{t}^{2}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}(Y_{i}-\hat{\mu}_{t}(x,a))^{2}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}},\ \hat{\mu}_{t}(x,a)=\frac{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}Y_{i}}{\sum\limits_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq t)}}.$ (9)
While the expressions may seem complicated, they are simply the sample mean and sample variance of the outcome, restricted to the subsample for group $x$ receiving treatment $a$. We remark that while the true underlying conditional means $\mu(x,a)$ and conditional variances $\sigma^{2}(x,a)$ do not depend on a specific stage $t$, their estimate are stage-specific. This is due to the sequential nature of the experiment: we always update previous estimates whenever new data become available.
@@
$(\tilde{e}_{t+1}(\cdot),\dots,\tilde{e}_{T}(\cdot))=\operatorname*{argmin}_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\ \hat{\mathsf{V}}_{t}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)),$ (10)
subject to the constraint that $\delta\leq e_{\ell}(\cdot)\leq 1-\delta$ for $\ell\geq t+1$. We then set $\hat{e}^{*}_{t+1}(1|\cdot)=1-\hat{e}^{*}_{t+1}(0|\cdot)=\tilde{e}_{t+1}(\cdot)$.
@@
$\hat{\mathsf{P}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)=$ $\sum_{\ell=1}^{t}\hat{r}_{\ell}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big{[}\hat{e}^{*}_{\ell}(1|x)\left(1-\hat{\mu}_{t}(x,1)\right)+\hat{e}^{*}_{\ell}(0|x)\left(1-\hat{\mu}_{t}(x,0)\right)\Big{]}$
$\quad+\sum_{\ell=t+1}^{T}r_{l}\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\Big{[}e_{\ell}(x)\left(1-\hat{\mu}_{t}(x,1)\right)+(1-e_{\ell}(x))\left(1-\hat{\mu}_{t}(x,0)\right)\Big{]}.$
@@
$\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}$ $\ \hat{\mathsf{P}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)$
$\text{s.t. }\hat{\mathsf{V}}_{t}\left(e_{t+1}(\cdot),\dots,e_{T}(\cdot)\right)+\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\big{(}\hat{\tau}_{t}(x)-\hat{\tau}_{t}\big{)}^{2}\leq C,$ (11)
where $\hat{\tau}_{t}(x)=\hat{\mu}_{t}(x,1)-\hat{\mu}_{t}(x,0)$ and $\hat{\tau}_{t}=\sum_{x\in\mathcal{X}}\hat{p}_{t}(x)\hat{\tau}_{t}(x)$. We then again set $\hat{e}^{*}_{t+1}(1|\cdot)=1-\hat{e}^{*}_{t+1}(0|\cdot)=\tilde{e}_{t+1}(\cdot)$.
As both optimization problems in (10) and (11) are fully forward-looking, our proposed design strategies aim to proactively adjust the allocation strategies in future stages, ensuring that the data collected accounts for the future impacts due to outcome delays. This approach contrasts with classical CARA designs that optimize trial objectives solely using historical information. At the end of the last stage $T$, the average treatment effect estimator is constructed with
@@
$\hat{\mathsf{V}}_{T}$ $=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\Bigg{[}\frac{\hat{\sigma}_{T}^{2}(x,1)}{\hat{\mathsf{e}}_{T}(x,1)}+\frac{\hat{\sigma}_{T}^{2}(x,0)}{\hat{\mathsf{e}}_{T}(x,0)}+\big{(}\hat{\tau}_{T}(x)-\hat{\tau}\big{)}^{2}\Bigg{]},\ \hat{\mathsf{e}}_{T}(x,a)=\frac{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x)}}.$
We summarize the analysis protocol below, where $z_{1-\frac{\alpha}{2}}$ denotes the $1-\frac{\alpha}{2}$ percentile of the standard normal distribution.
Protocol for interim analysis and statistical inference
Initial stage 1:
   Enroll $n_{1}$ participants, and assign treatments with $\hat{e}_{1}^{*}(1|x)=\frac{1}{2}$.
Interim analysis after stage $t$ has concluded:
   Available data: $(S_{i},X_{i},A_{i},D_{i},Y_{i})$ for $i=1,2,\dots,N_{t}$;
   Following (7) and (8), construct the delay distribution estimate $\hat{\rho}_{t}(d|x,a)$;
   Following (9), construct the outcome distribution estimates $\hat{\sigma}_{t}(x,a)$ and $\hat{\mu}_{t}(x,a)$;
   From either (10) or (11), solve the allocation for stage $t+1$: $\hat{e}_{t+1}^{*}(1|x)$.
Treatment assignment for stage $t+1$:
   Enroll $n_{t+1}$ participants, and assign treatments with $\hat{e}_{t+1}^{*}(1|x)$.
Statistical inference after the final stage $T$ has concluded:
   Available data: $(S_{i},X_{i},A_{i},D_{i},Y_{i})$ for $i=1,2,\dots,N$;
   Construct the treatment effect estimator $\hat{\tau}_{T}$ and the variance estimator $\hat{\mathsf{V}}_{T}$;
   Report the $(1-\alpha)\%$ confidence interval: $\hat{\tau}_{T}\pm z_{1-\frac{\alpha}{2}}\sqrt{\hat{\mathsf{V}}_{T}/N}$.
In the following sections, we will provide theoretical guarantees and simulation evidence to demonstrate the advantages of our proposed fully forward-looking CARA designs in the presence of delayed outcomes.